OpenStudy (anonymous):

Which of the following would you use as the percent equation to determine that 40 is 70% of what number?

OpenStudy (whpalmer4):

\[40 = 70\% * x = \frac{70}{100}*x\]\[x= \frac{40*100}{70}\]

OpenStudy (jhannybean):

\[\large \frac{70}{100}=\frac{40}{x}\] this tells you 70 as a percentage, and you're taking 70 % of the 40 you're given.\[70x = 4000 \ \therefore \ x= \frac{4000}{70}\]

OpenStudy (anonymous):

40= 0.7 x a